# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

### v1

@@ 8. Conclusions and Future Work
While predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation. This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation methods, and imputation techniques that are grounded in the lifecycle of data collection, processing and use. That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the benchmark implementation, which include a broad description of the missing value imputation techniques used in our study and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair benchmark (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix C.1 presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4. Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 6. Specifically, Appendix D.2 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.3 explores further experimental results on model performance under missingness shift with variable train and test error rates.
Appendix A Additional details on benchmark implementation
@@ A.1. Missing Value Imputation Techniques
In our benchmark, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021). In each category, we selected one or two SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
@@ A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Prior to constructing our benchmark, we conducted a thorough analysis of the strengths and weaknesses of existing state-of-the-art benchmarks and MVI techniques by reviewing their codebases. Unfortunately, we identified several methodological flaws and code bugs in some of them. Consequently, alongside creating our software novelty in the benchmark, we dedicated considerable effort to correcting these bugs in the null imputers and enhancing existing benchmarking approaches to ensure the development of a truly fair benchmark.
Although our benchmark pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the benchmark with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Consequently, the authors chose to use these imputers without modification. In contrast, our benchmark aimed to ensure a fair comparison of MVI techniques as they would be used in production systems. To achieve this, we fixed and extended existing MVI techniques, introducing another software innovation in our benchmark.
In our study, we standardized all 8 MVI techniques discussed in Section 3.2 to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameters tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below are the enhancements we implemented for MVI techniques used in our benchmark:
  • clustering: We implemented this method ourselves to meet our imputation requirements, as existing solutions do not support null imputation based on clustering for datasets containing both numerical and categorical features. For this, we used the k-prototypes as a distance function in the imputer. [^8] https://github.com/nicodv/kmodes
@@
In summary, our benchmark offers users not only a robust evaluation framework but also incorporates the refined MVI techniques developed as part of our work.
@@ B.1. Computing Infrastructure
Our large-scale experimental study incorporated 8 null imputation techniques, 5 model types, 6 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 20,952 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our benchmark controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
@@ B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 4-9 provide the demographic composition of all six datasets, expressed through the proportions and base rates of each protected group. Figures 13-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 10-14 detail the missingness scenarios for the folk, bank, heart, law-school, and german datasets, respectively.
@@
Table 5. Table 5: Proportions and Base Rates for folk.
@@
Table 7. Table 7: Proportions and Base Rates for diabetes.
overall gender_priv gender_dis
Proportions 1.0 0.621 0.379
Base Rates 0.291 0.272 0.321
@@
Table 10. Table 10: Missingness scenarios for an error rate of 30% for folk. WKHP, AGEP are numerical columns; SCHL, MAR are categorical columns.
@@
Figure 13. EDA for designing missingness scenarios in folk.
@@ Appendix C Additional Experimental Results for Single-mechanism Missingness
C.1. Imputation Quality and Fairness
To better understand the impact of different MVI techniques on model performance, we measure the imputation quality of each imputer before training ML models on the imputed training set. The quality of imputation is evaluated across three dimensions: correctness (using F1 for categorical features and RMSE for numerical features), distributional alignment (using KL divergence of imputed values), and fairness (measuring differences in F1, RMSE, and KL divergence between disadvantaged and privileged groups). The higher values of F1 are the better, the lower values of RMSE and KL divergence are the better, and values of imputation fairness closer to zero are the better.
To reduce the number of plots, we selected one categorical and one numerical features for each dataset based on their Spearman correlation with the target. However, our analysis takes into account the imputation quality measured on all features with nulls. Importantly, these figures do not include boostclean since this is a joint cleaning and model training method that does not provide imputed values, which we could use to measure imputation quality. The x-axis labels include the names of the corresponding features from the dataset for which the imputation quality metric was measured.
Correctness of imputation. Figures 18 and 19 demonstrate the correctness of imputation of all MVI techniques, except boostclean, for all 6 datasets. From Figure 18, we can notice that one of the ML and DL-based imputers (miss-forest, auto-ml, datawig) is the best in terms of F1 for categorical columns under MCAR and MAR missingness mechanisms. Additionally, Figure 19 shows that all the three ML and DL-based imputers are dominant in terms of RMSE for numerical columns compared to statistical MVI techniques. However, these significant differences in F1 and RMSE do not translate into substantial differences in downstream model performance, as measured by F1 in Section 4.1. This observation further supports our statement made in Section 5.4 that ”imputation quality metrics are not very meaningful because a bad data imputation method can be overcome by a strong learner, and so bad performance on imputation quality does not necessarily mean than the MVI will lead to poor predictive performance.”
Distributional alignment of imputation. Figures 20 and 22 illustrate the distributional alignment of imputation of all MVI techniques except boostclean, for all 6 datasets. The plots clearly demonstrate that ML and DL-based imputers, except clustering, outperform statistical approaches in terms of distributional alignment across all missingness mechanisms. In Section 5.4, we highlight scenarios where distributional metrics serve as more reliable indicators than discrepancy metrics. This is because the distributional alignment of imputed features can significantly influence model training outcomes, emphasizing that assessing only the correctness of imputation is often insufficient.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 18. F1 of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 19. RMSE of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 20. KL divergence of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 21. F1 Difference of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 22. KL divergence of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 23. RMSE Difference of different imputation strategies for the most correlated numerical column.
Fairness of imputation. Figures 21, 23, 24, and 25 show the fairness of imputation of all MVI techniques, except boostclean, for all 6 datasets. We report fairness metrics based on an intersectional group (i.e., sex&race) for datasets with two sensitive attributes, and a singular group (i.e., age) for datasets with one sensitive attribute. From Figures 21 and 23, we can notice that statistical MVI techniques are prone to worsen imputation fairness rather than improve it under all missingness mechanisms. Further, comparing these imputation fairness differences with the model fairness differences discussed in Section 4, we can state that good imputation fairness does not guarantee good model fairness.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 24. KL Divergence Difference of different imputation strategies (colors in the legend) for the most correlated categorical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 25. KL Divergence Difference of different imputation strategies (colors in the legend) for the most correlated numerical column with the target per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
C.2. Additional Accuracy and Fairness Metrics of Predictive Model
@@
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular fairness metrics of downstream model performance. Figures 26, 27, and 28 illustrate Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
@@
Overall, the plots support the assertion made in Section 4.2 that the impact of MVI on fairness (error-disparity) is highly dataset-specific and is most strongly influenced by the baseline disparity of the model trained on clean (complete, no missingness) data. This implies that it is impossible to determine in advance whether data cleaning will worsen fairness.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 26. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 27. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 28. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a clean baseline.

### v2

@@ 7. Conclusions, Limitations and Future Work
It is interesting future work to create research datasets with naturally-occurring missingness shifts. Further, while predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation. This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation procedures, and imputation techniques that are grounded in the lifecycle of data collection, processing and use. That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix LABEL:apdx:single-mechanism-imp-quality presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.1 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4. Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates.
Appendix A Additional details on the evaluation suite implementation
@@ A.1. Missing Value Imputation Techniques
In our evaluation suite, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021). In each category, we selected one or two SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
@@ A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Prior to constructing our evaluation suite, we conducted a thorough analysis of the strengths and weaknesses of existing state-of-the-art benchmarks and MVI techniques by reviewing their codebases. Unfortunately, we identified several methodological flaws and code bugs in some of them. Consequently, alongside creating our software novelty in the evaluation suite, we dedicated considerable effort to correcting these bugs in the null imputers and enhancing existing benchmarking approaches to ensure the development of a truly fair evaluation suite.
Although our evaluation suite pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the evaluation suite with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Presumably, the authors chose to use these imputers without modification. In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized all 8 MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below is a summary of the enhancements we implemented for MVI techniques used in our evaluation suite:
  • clustering: We implemented this method ourselves to meet our imputation requirements, as existing solutions do not support null imputation based on clustering for datasets containing missingness in both numerical and categorical features at the same time. For this, we used the k-prototypes as a distance function in the imputer. [^9] https://github.com/nicodv/kmodes
@@
In summary, our evaluation suite offers users not only a robust evaluation framework but also incorporates fixes to the MVI implementations previously available.
@@ B.1. Computing Infrastructure
Our large-scale experimental study incorporated 8 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 23,940 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our evaluation suite controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
@@ B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 4-10 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group. Figures 12-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 11-16 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
Table 4. Table 4: Proportions and Base Rates for diabetes.
overall gender_priv gender_dis
Proportions 1.0 0.621 0.379
Base Rates 0.291 0.272 0.321
@@
Table 6. Table 6: Proportions and Base Rates for folk-income.
@@
Table 10. Table 10: Proportions and Base Rates for folk-employment.
Proportions 1.0 0.490 0.510 0.625 0.375 0.806 0.194
Base Rates 0.57 0.617 0.525 0.563 0.582 0.577 0.541
Table 11. Table 11: Missingness scenarios for an error rate of 30% for folk-employment. AGEP is a numerical column; DIS, MIL, SCHL are categorical columns.
MCAR DIS, MIL, AGEP, SCHL N/A 0.3 0.3
MAR MIL, AGEP SEX 0.2 (female) 0.1 (male)
DIS, SCHL RAC1P 0.2 (non-white) 0.1 (white)
MNAR DIS DIS 0.25 (with disability) 0.05 (without disability)
[…]
AGEP AGEP 0.25 (¿ 50) 0.05 ($\leq 50$)
SCHL SCHL 0.25 (¡ 21) 0.05 ($\geq 21$)
(a) Correlation with label
(b) Feature importance
Figure 12. EDA for designing missingness scenarios in folk-employment.
Table 12. Table 12: Missingness scenarios for an error rate of 30% for folk-income. WKHP, AGEP are numerical columns; SCHL, MAR are categorical columns.
@@
Figure 13. EDA for designing missingness scenarios in folk-income.
@@ Appendix C Additional Experimental Results for Single-mechanism Missingness
C.1. Additional Accuracy and Fairness Metrics of Predictive Model
@@
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance. Figures 18, 19, 21, and 20 illustrate Accuracy, Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@
Overall, the plots support the assertion made in Section 4.2 that the impact of MVI on fairness (error-disparity) is most strongly influenced by the baseline disparity of the model trained on clean (complete, no missingness) data. This implies that it is hard to determine in advance whether data cleaning will worsen fairness.
@@
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 19. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 20. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 21. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.